OpenStudy (hihi67):

the square of a number is 8 smaller than the product of the number and its successor. what is the number, please show in an equation too.

OpenStudy (hihi67):

do you know what it is?

OpenStudy (anonymous):

try starting by using x as the number they are describing.

OpenStudy (hihi67):

i know. i did x^2

OpenStudy (hihi67):

okay. and the answer is 8, i just don't have an equation.

OpenStudy (anonymous):

\[x ^{2}=(x)(x+1)-8\]

OpenStudy (hihi67):

okay. i had \[x^2 = 8-x*(x+1)\]

OpenStudy (hihi67):

does that work?

OpenStudy (anonymous):

no

OpenStudy (hihi67):

why not?

OpenStudy (anonymous):

the number squared is 8 LESS than the the product of the two.

OpenStudy (anonymous):

pg stated it correctly \[x^{2}=x ^{2}+x-8; so 0=x-8 ; so x=8\]